alldc(1) teTeX alldc(1) NAME alldc−forcethemostimportantec−fontstobecalculated SYNOPSIS alldc DESCRIPTION alldcforcesalargenumberofec-styleFontstobecalculatedaspixelfiles.Thisisdonethrough runningdvips(1)overacertaintest−file.Therefore,thefontsarecreatedintheresolutionneeded bydvips. alldcdoesnotrecalculateexistingfonts(aslongastheMetafontmodedoesnotchange). SEEALSO allcm(1),allneeded(1),fontimport(1),fontexport(1). AUTHOR ThomasEsser(te@informatik.uni−hannover.de) teTeX 11/94 1